(x^2+1) y''+xy'-xy=0 in series solution
1
Expert's answer
2020-04-21T16:42:59-0400

Given

"(x^2+1) y''+xy'-xy=0"

Let

"\begin{aligned}
 y&=\sum_{n=0}^{\infty} a_{n} x^{n}\\
 y'&=\sum_{n=1}^{\infty} na_{n} x^{n-1}\\
 y''&=\sum_{n=2}^{\infty} n(n-1)a_{n} x^{n-2}
\end{aligned}"

Then

"\begin{aligned}
 (x^2+1)\sum_{n=2}^{\infty} n(n-1) a_{n} x^{n-2}+x \sum_{n=1}^{\infty} na_{n} x^{n-1}-x\sum_{n=0}^{\infty} a_{n} x^{n}&=0\\
 \sum_{n=2}^{\infty} n(n-1) a_{n} x^{n}+\sum_{n=2}^{\infty} n(n-1) a_{n} x^{n-2} + \sum_{n=1}^{\infty} na_{n} x^{n}-\sum_{n=0}^{\infty} a_{n} x^{n+1}&=0\\
 \sum_{k=2}^{\infty} k(k-1) a_{k} x^{k}+\sum_{k=4}^{\infty}( k-2)(k-3) a_{k-2} x^{k} + \sum_{k=1}^{\infty} ka_{k} x^{k}-\sum_{k=1}^{\infty} a_{k-1} x^{k}&=0\\
2a_2x^2+6a_3x^3+\sum_{k=4}^{\infty} k(k-1) a_{k} x^{k}+ \sum_{k=4}^{\infty}( k-2)(k-3) a_{k-2} x^{k} \\
 +a_1x+2a_2x^2+3a_3x^3 + \sum_{k=4}^{\infty} ka_{k} x^{k}-a_0x-a_1x^2-a_2x^3-\sum_{k=4}^{\infty} a_{k-1} x^{k}&=0\\
(a_1-a_0)x+(4a_2-a_1)x^2+(9a_3-a_2)x^3\\
+\sum_{k=4}^{\infty}[ k(k-1) a_{k}+(k-2)(k-3)a_{k-2}+ka_k-a_{k-1} ]x^k&=0

\end{aligned}"

Then we get

"\begin{aligned}
a_1-a_0&=0\ \ \to a_1=a_0\\
4a_2-a_1&=0\ \ \to a_2= \frac{1}{4}a_0\\
9a_3-a_2&=0\ \ \to a_3=\frac{1}{36}a_0\\
k^2a_k&=(k-2)(k-3)a_{k-2}-a_{k-1}\\
a_{k}&=\frac{(k-2)(k-3)a_{k-2}-a_{k-1}}{k^2},\ \ \ \ \ k\geq 4
\end{aligned}"

Hence

"\begin{aligned}

\end{aligned}" "\begin{aligned}
a_4&=\frac{2a_2-a_3}{16}=\frac{17}{576}a_0\\
a_5&=\frac{2a_2-a_3}{16}=\frac{79}{14400}a_0\\
\vdots
\end{aligned}"

Then we can write the solution as

"\begin{aligned}
y&=\sum_{n=0}^{\infty}a_nx^n\\
&=a_0+a_1x+a_2x^2+a_3x^3+a_4x^4+\cdots \\
&=a_0+a_0x+\frac{1}{4}a_0x^2 +\frac{1}{36}a_0x^3+\frac{17}{576}a_0x^4+\frac{79}{14400}a_0x^5+\cdots 
\end{aligned}"